& Quizzes My Open Math Assignments > Assessment plots Due in 14 hours, 3 minutes 6 78 910 1 12 13 14 15 data Based on the boxplot above, identify the 5 number summary (min, Q1, median, Q3, max) The Inter Quatile Range QR)-03-Q1 Hint: Help C Video Textbook Points possible: 1 This is attempt 2 of 3. Submit
Answer #1

# Based on the box plot above ,the five number summary is

min=6

Q1 =7

Median = 13

Q3= 14

max= 15

# The inter-quartile range IQR = Q3 - Q1 = 14 - 7= 7

IQR = 7